# thanks: Author to whom correspondence should be addressed

### v1

@@ Abstract
Understanding transport phenomena through porous media is essential for applications ranging from water treatment systems to heat pipes. In many of these systems, packed-bed reactors (PBRs) are crucial components, and understanding and quantifying the pressure drop due to flow through the PBR is critical to effective operation. Recent experiments conducted by NASA measured the pressure drop due to gas-liquid flow through a PBR under microgravity conditions. Based on these experiments, we develop correlations for the interphase drag in a two-fluid model (TFM). Specifically, two closure relations are needed for the TFM: the liquid-solid $f_{ls}$ and gas-liquid $f_{gl}$ interphase force. We use an Ergun-type closure for $f_{ls}$. Then, under a 1D flow assumption, the TFM equations are rewritten with $f_{gl}$ as the only unknown. We employ data-driven calculations to determine $f_{gl}$, which we correlate (via composite fits) as a function of the liquid and gas Reynolds numbers, $Re_{l}$ and $Re_{g}$, respectively, and the Suratman number $Su_{l}$. To validate the proposed $f_{gl}(Re_{l},Re_{g},Su_{l})$ closure, we perform two-dimensional (2D) transient, multiphase computational fluid dynamics (CFD) simulations at low $Re_{l}$ and $Re_{g}$ (laminar flow) in ANSYS Fluent employing an Euler-Euler formulation. We find good agreement between the CFD simulations based on the proposed $f_{gl}$ closure and the experimental data.
@@ 1 Introduction
Multiphase flows through porous media (Adler and Brenner, 1988; Wooding and Morel-Seytoux, 1976) find applications across various systems, from heat pipes to volatile removal assembly and hydrogenation plants, to name a few (Azarpour, Rezaei, and Zendehboudi, 2021). One such problem is gas-liquid (two-phase) flow through a packed-bed reactor (PBR). A typical PBR consists of spherical catalyst particles packed within a cylindrical (or cuboidal) pipe (Ranade, Chaudhari, and Gunjal, 2011). In the presence of gravity, the gas and liquid phases arrange themselves in the trickle flow regime (the liquid phase trickles over the packing in a gas continuous medium), and the corresponding pressure drops and trickle-to-pulse flow regime transitions have been studied extensively (see, for instance, Refs. Wang, Chen, and Larachi, 2013; Attou, Boyer, and Ferschneider, 1999; Holub, Duduković, and Ramachandran, 1992, 1993 and those therein). On the other hand, in the absence of gravity (or under microgravity conditions), the capillary force becomes important. Bubble flow (gas droplets in liquid continuous medium) and pulse flow (alternating regions of gas-rich and liquid-rich) regimes are observed in two-phase flow through PBRs at low-gravity conditions (Motil et al., 2021; Motil, Balakotaiah, and Kamotani, 2003; Taghavi et al., 2022). However, determining the pressure drop across the PBR at microgravity conditions remains an active research problem (Balasubramaniam et al., 2006). Recently, progress has been made via new experimental and theoretical efforts.
@@
Specifically, Motil, Balakotaiah, and Kamotani (2003) reported on parabolic flight experiments of gas-liquid flow through a PBR. By developing flow regime maps, they demonstrated that different flow physics compete in microgravity conditions compared to terrestrial gravity conditions. In addition, a new correlation known as the “modified Ergun equation” was proposed to predict the frictional pressure drop for gas-liquid flows through PBRs at microgravity conditions. The duration of low-gravity conditions on parabolic flights is short. To remedy this shortcoming and gather data on longer-duration PBR flows, further experiments were conducted aboard the International Space Station (ISS)(Motil et al., 2021; Taghavi et al., 2022). Specifically, the wettability and hysteresis effects were studied in the viscous-capillary regime (low Reynolds number flow) by measuring pressure drops and bubble-to-pulse flow regime transitions for Teflon (nonwetting) and glass (wetting) packing materials (Motil et al., 2021). These experiments suggested that contributions from capillarity dominate the pressure drop for wetting packing, whereas contributions from viscous drag dominate the pressure drop for the nonwetting cases. Further, the effect of pre-flow (i.e., gas-flush versus liquid-flush) through a PBR on the pressure drop was investigated (Taghavi et al., 2022). Higher gas holdup and pressure drops were observed for liquid pre-flow conditions compared to gas pre-flow conditions. More recently, a novel technique to characterize flow regime transitions via the slope of the pressure drop versus flow rate curve (from the ISS data) demonstrated the dominance of capillarity at low flow rates and the dependence of the gas holdup on pre-flow conditions (Taghavi et al., 2023).
@@
One-dimensional (1D) two-fluid models (TFMs) have been developed to understand these experimental observations from a fundamental (continuum) point of view (Gidaspow, 1994). Assuming the flow to be 1D, uniform in the cross-section, and at steady-state, the full TFM equations can be simplified so that the unknowns are the axial pressure drop, the liquid holdup, and the interphase drag force (Zhang et al., 2017; Salgi and Balakotaiah, 2014, 2017). Following this approach, Zhang et al. (2017) performed experiments and developed a 1D TFM at low Bond number $Bo$ (the ratio of gravitational to surface tension forces), within which the trickle flow regime is dominant. They incorporated Ergun-type correlations in the three interphase drag forces (i.e., the liquid-solid $f_{ls}$, gas-liquid $f_{gl}$, and gas-solid $f_{gs}$ ones) with appropriate relative velocities (Attou, Boyer, and Ferschneider, 1999; Holub, Duduković, and Ramachandran, 1992, 1993). At low $Bo$, capillarity implies a pressure jump (drop) across a gas-liquid interface: $p_{c}=p_{g}-p_{l}\neq 0$, where $p$ is pressure and the subscripts “$c$”, “$g$”, and “$l$” stand for “capillary,” “gas”, and “liquid,” respectively. For example, Zhang et al. (2017) employed the correlation from Attou and Ferschneider (2000) to model $p_{c}$ as a function of surface tension, particle diameter, gas holdup, and the ratio of gas and liquid densities. Next, the shared pressure in the 1D gas and liquid phases’ momentum equations was eliminated to obtain a single momentum equation, with the liquid holdup being the only unknown. Finally, the pressure drop across the PBR was calculated by substituting the predicted liquid holdups, which were determined for different gas and liquid flow rates, obtaining a good agreement between the 1D TFM predictions and the experiments.
@@
Meanwhile, Salgi and Balakotaiah (2014) used a modified Ergun correlation to calculate the pressure drop under microgravity conditions (Motil, Balakotaiah, and Kamotani, 2003). They treated the liquid holdup and $f_{gl}$ as unknown in their model. They used a so-called “dynamic interaction term” from Motil, Balakotaiah, and Kamotani (2003) to improve the fit to the experimental pressure drops. Further, a capillary pressure drop correlation, derived under terrestrial gravity conditions, was incorporated because there are no acceptable capillary pressure drop correlations for microgravity conditions. A linear stability analysis was performed to obtain $f_{gl}$ and liquid holdups at different gas and liquid flow rates. Similarly, Salgi and Balakotaiah (2017) developed experimentally-based correlations for $f_{gl}$ from pulse velocity and frequency data to predict pressure drops through PBRs at both microgravity and terrestrial gravity conditions.
@@
In these prior 1D TFM studies, model calibration (of one or more fitting parameters) is key to accurately predicting pressure drops. Further, as Motil et al. (2021) stated in their paper, “Ergun type expressions for $f_{gl}$ cannot be justified in the dispersed bubble flow regime.” However, it is the dispersed bubble and pulse flow regimes that are observed at microgravity conditions (Motil et al., 2021; Motil, Balakotaiah, and Kamotani, 2003), which we seek to model in this work. Therefore, the goal of the present work is to perform a data-driven calculation of the drag coefficient (fitting coefficient of our model) to obtain composite fits that reveal a novel correlation for $f_{gl}$ at microgravity conditions. To the best of our knowledge, this is one of the first gas-liquid interphase drag force correlations for flows through PBRs at microgravity conditions.
@@
In addition to theoretical efforts, computational fluid dynamics (CFD) has been widely used to study resolved two-phase flows through PBRs. Specifically, an Euler-Euler TFM approach is used to predict pressure drop and liquid holdup within PBRs. Previous CFD studies(Atta, Roy, and Nigam, 2007b, a; Lopes and Quinta-Ferreira, 2008; Lu et al., 2018; Gunjal et al., 2005) considered an “empty” channel and incorporated the relevant closure relations for interphase drag terms (such as $f_{gl}$, $f_{ls}$, and $f_{gs}$) as momentum source terms, which implicitly accounts for the porous medium. In this modeling approach, one does not have to resolve the complex flow domain of interstitial void spaces to extract relevant flow parameters (Nagrani, Marconnet, and Christov, 2024). In the present study, we follow this approach and incorporate the proposed $f_{gl}$ closure relations (developed from the 1D TFM model) as momentum source terms in CFD simulation. Then, we compare the pressure drop from these CFD simulations with the experimental pressure drop measured by NASA’s Packed Bed Reactor Experiment (PBRE), provided as data in NASA’s Physical Sciences Informatics (PSI) system (Motil, Mitchell, and Balakotaiah, 2020). This comparison serves as an example demonstration of the utility of the proposed $f_{gl}$ correlations.
@@ 2 Mathematical model
Our goal is to obtain a novel closure relation for interphase gas-liquid drag force $f_{gl}$ in a TFM, based on microgravity experiments (Motil, Mitchell, and Balakotaiah, 2020). The idea of a TFM is to treat the carrier and dispersed phases of a two-phase flow as interpenetrating continua, each having its own thermophysical properties and obeying suitable balance laws (Gidaspow, 1994). The two continua “interact” through the interphase drag force, denoted $f$. Therefore, the interphase drag force couples the momentum equations of the two phases. Throughout, it should be understood that $f$ usually represents a body force density (i.e., a force per unit volume), but we refer to it as “force” for clarity and simplicity.
In this section, we present our proposed procedure for obtaining $f_{gl}$. Specifically, in Sec. 2.1, we start from the basic mass and momentum balances under the TFM. In Sec. 2.2, we employ experimental data from NASA’s PBRE (Motil, Mitchell, and Balakotaiah, 2020) to perform a data-driven inference and calibration of a model for $f_{gl}$. In Sec. 2.3, we demonstrate the self-consistency of the modeling approach by comparing the predicted pressure drops with the original experimental data.
@@ 2.1 Governing equations
The governing equations of gas-liquid flow through a PBR are those of conservation of mass and conservation of momentum. We assume the flow is predominantly 1D, with gradients along the axial (i.e., flow) direction of PBR being of interest. Under this assumption, the conservation of mass for liquid and gas phases, respectively, can then be simplified to:
@@
In a 1D TFM model of two-phase flow through a PBR, it is generally assumed (an approximation is made) that the mesoscale velocities are $v_{g}=v_{gs}/(\epsilon\phi_{g})$ and $v_{l}=v_{ls}/(\epsilon\phi_{l})$, where $v_{gs}$ and $v_{ls}$ are the corresponding gas or liquid superficial velocities, and $\epsilon$ is the PBR void fraction (porosity) (Salgi and Balakotaiah, 2014, 2017; Zhang et al., 2017). Next, the 1D conservation of momentum equations for the liquid and gas phases (based on the mesoscale velocities), respectively, are
@@
To close the TFM (Eqs. (1) and (2)), we need to specify the interphase momentum transfer terms, which are represented by the forces $f_{g}$ and $f_{l}$. Note that the expressions for $f_{g}$ and $f_{l}$ can each have a number of additive contributions, depending on the physics at hand. To arrive at expressions for $f_{g}$ and $f_{l}$, we assume that the interaction between the gas phase and the solid packing is negligible, i.e., $f_{gs}=0$, because the liquid wets the solid packing (Salgi and Balakotaiah, 2014, 2017). Then, we write the phase-specific forces in terms of the interphase ones as $f_{l}=f_{gl}-f_{ls}$ and $f_{g}=-f_{gl}$. We remind the reader that $f_{ls}$ is the interphase drag force between the liquid and solid phases, and $f_{gl}$ is the interphase drag force between the gas and liquid phases.
Next, it is standard to assume fully developed flow at steady state (Salgi and Balakotaiah, 2014, 2017; Zhang et al., 2017). Further, a shared pressure exists between the gas and liquid phases: $p_{g}=p_{l}=p$. Incorporating these assumptions into Eqs. (2), we have
@@ 2.2.1 Two-phase Ergun correlation for the liquid-solid drag
Previous studies(Salgi and Balakotaiah, 2014, 2017; Zhang et al., 2017) used an Ergun-type (Ergun and Orning, 1949; Ergun, 1952) equation to determine $f_{ls}$. An Ergun-type equation consists of additive viscous and inertial contributions to the pressure gradient across a PBR, but the relation can also be used to express the drag. Specifically, Salgi and Balakotaiah (2014, 2017) adapted the Ergun-type equation:
@@
where $d_{p}$ is the diameter of the spherical particles comprising the packing. Here, $f_{ls}=A_{ls}$ is a single-phase Ergun-type (Ergun and Orning, 1949; Ergun, 1952) equation for the liquid phase, which has been adjusted by the denominator $\phi_{l}^{2}$ to account for the presence of a gas phase reducing the area available for the flow of the liquid phase. In the present study, we use Eq. (4) to model the drag between the liquid and solid phases where the coefficients $180$ and $1.8$ for vicious and inertia components respectively are from Motil et al. (2021).
@@ 2.2.2 Data-driven calibration of the gas-liquid drag
Next, we propose a novel way of obtaining a closure relation (i.e., a correlation) for $f_{gl}$ from experimental PBR pressure drop data obtained from NASA’s PBRE, which we accessed via NASA’s PSI (Motil, Mitchell, and Balakotaiah, 2020). As indicated in the works of Salgi and Balakotaiah (2014), bubble and pulse flow regimes are predominant at microgravity conditions, and the trickle flow regime does not occur. Hence, one cannot use an Ergun-type correlation for $f_{gl}$, with a relative velocity because that is valid only in the trickle flow regime (Attou, Boyer, and Ferschneider, 1999; Holub, Duduković, and Ramachandran, 1992, 1993). To circumvent this issue, we use a traditional drag force formulation: $f_{gl}=\beta(v_{g}-v_{l})$, where $\beta$ is the drag coefficient, and $v_{g}-v_{l}$ is the relative velocity. Using this formulation, the shared pressure gradient can be eliminated between Eqs. (3a) and (3b), while $f_{ls}$ is eliminated via Eq. (4), to obtain a single governing equation:
@@
To justify the shared pressure assumption in the present work, we note the absence of an acceptable correlation for the capillary pressure drop at microgravity conditions (Salgi and Balakotaiah, 2014). If $p_{g}-p_{l}=p_{c}\neq 0$ (i.e., we cannot let $p=p_{g}=p_{l}$), then a term proportional to $p_{c}/d_{p}$ would arise in Eq. (5), which is a priori unknown. In the simplest case of a static gas bubble, the Young–Laplace law governs the pressure jump, $p_{c}$, across the gas-liquid interface due to surface tension. However, the gas-liquid interface continuously evolves throughout the PBR and hence cannot be modeled using just the Young–Laplace law (Taghavi et al., 2023). Nevertheless, we can still implicitly account for capillarity through the dependence of $f_{gl}$ (or, $\beta$) on an appropriate Suratman number (a combined measure of the relative importance of surface tension to both inertial and viscous forces in the flow).
@@
To quantify this dependence, we performed a set of linear curve fits on the $f_{gl}$ versus $Re^{*}_{gs}$ data at each $Re^{*}_{ls}$ to obtain correlations of the form: $f_{gl}=a+bRe^{*}_{gs}$, where $a$ and $b$ are, respectively, the intercept and slope of the curvefit at fixed value of $Re^{*}_{ls}$. To combine these multiple correlations into a single correlation, we perform a composite linear fit to the coefficients $a$ and $b$ (obtained across all $Re^{*}_{ls}$ values) with respect to $Re^{*2}_{ls}/Su_{l}$, where $Su_{l}=\rho_{l}d_{p}\sigma/\mu_{l}^{2}$ is the Suratman number, as shown in Fig. 1(b) and Fig. 1(c) of the forms: $a(\xi)=\alpha+\beta\xi$ and $b(\xi)=\gamma+\delta\xi$. We introduced $Su_{l}$ here to implicitly account for surface tension (capillarity effects) between the gas and liquid phases.
@@
$f_{gl}^{\text{Teflon}}$ $=291.6+4.22\times 10^{4}\frac{Re^{*2}_{ls}}{Su_{l}}$
$\quad+36.2Re^{*}_{gs}+4517\frac{Re^{*2}_{ls}Re^{*}_{gs}}{Su_{l}},$ (6a)
$f_{gl}^{\text{glass}}$ $=330.8+9.92\times 10^{4}\frac{Re^{*2}_{ls}}{Su_{l}}+27.2Re^{*}_{gs}$
$\quad+1.466\times 10^{4}\frac{Re^{*2}_{ls}Re^{*}_{gs}}{Su_{l}}.$ (6b)
Finally, if one were to have accepted correlations for the capillary pressure in a two-phase flow in microgravity conditions, it can be added back into Eq. (5). In that case, the fitting process can be repeated, and the curve-fit coefficients would change. Thus, if one wanted to explicitly incorporate capillarity effects into this approach, one would obtain different coefficient values and dependencies, but the methodology remains the same.
Figure 3: Verification of the proposed $f_{gl}$ correlations in Eq. (6) (and hence the approach) by comparing the model and experiment’s $-dp/dz$ variation with $Re^{*}_{gs}$ at different $Re^{*}_{ls}$ for (a) Teflon packing and (b) glass packing PBRE data. As expected, the model predictions ($\bigtriangleup$) at discrete data points coincide exactly with the values from experiments ($\times$). The arrow indicates the direction of increasing $Re^{*}_{ls}$. The dashed lines are a guide to the eye and do not represent any trends.

### v2

# Correlations for the interphase drag in the two-fluid model of gas-liquid flows through packed-bed reactors

@@ Abstract
Experiments conducted by NASA measured the pressure drop due to gas-liquid flow through a packed-bed reactor under microgravity conditions. From these experiments, we develop correlations for the gas-liquid $f_{gl}$ interphase drag in a two-fluid model (TFM). We use an Ergun-type closure for liquid-solid drag. Then, under a 1D flow assumption, $f_{gl}$ is the only unknown in the TFM. Using a data-driven approach, we determine $f_{gl}$ and correlate it (via composite fits) with the liquid and gas Reynolds numbers, $Re_{l}$ and $Re_{g}$, respectively, and the Suratman number $Su_{l}$. To validate the proposed $f_{gl}(Re_{l},Re_{g},Su_{l})$ closure, we perform two-dimensional transient simulations at microgravity conditions using ANSYS Fluent and employing an Euler-Euler formulation. We find good agreement between the simulations based on the proposed $f_{gl}$ closure and the experimental data.
keywords:
American Chemical Society,
[…]
@@ 1 Introduction
Multiphase flows through porous media 1, 2 find applications across various systems, from heat pipes to volatile removal assembly and hydrogenation plants, to name a few 3. One such problem is gas-liquid (two-phase) flow through a packed-bed reactor (PBR). A typical PBR consists of spherical catalyst particles packed within a cylindrical (or cuboidal) pipe 4. In the presence of gravity, the gas and liquid phases arrange themselves in the trickle flow regime (the liquid phase trickles over the packing in a gas continuous medium), and the corresponding pressure drops and trickle-to-pulse flow regime transitions have been studied extensively (see, for instance, Refs. 5, 6, 7, 8 and those therein). On the other hand, in the absence of gravity (or under microgravity conditions), the capillary force becomes important. Bubble flow (gas droplets in liquid continuous medium) and pulse flow (alternating regions of gas-rich and liquid-rich) regimes are observed in two-phase flow through PBRs at low-gravity conditions 9, 10, 11. However, determining the pressure drop across the PBR at microgravity conditions remains an active research problem 12. Recently, progress has been made via experimental and modeling efforts.
@@
Specifically, Motil et al. 10 reported on parabolic flight experiments of gas-liquid flow through a PBR. By developing flow regime maps, they demonstrated that different flow physics compete in microgravity conditions compared to terrestrial gravity conditions. In addition, a correlation known as the “modified Ergun equation” was proposed to predict the frictional pressure drop for gas-liquid flows through PBRs at microgravity conditions. The duration of low-gravity conditions on parabolic flights is short. To remedy this shortcoming and gather data on longer-duration PBR flows, further experiments were conducted aboard the International Space Station (ISS)9, 11. Specifically, the wettability and hysteresis effects were studied in the viscous-capillary regime (low Reynolds number flow) by measuring pressure drops and bubble-to-pulse flow regime transitions for Teflon (nonwetting) and glass (wetting) packing materials 9. These experiments suggested that contributions from capillarity dominate the pressure drop for wetting packing, whereas contributions from viscous drag dominate the pressure drop for the nonwetting cases. Further, the effect of pre-flow (i.e., gas-flush versus liquid-flush) through a PBR on the pressure drop was investigated 11. Higher gas holdup and pressure drops were observed for liquid pre-flow conditions compared to gas pre-flow conditions. More recently, a technique to characterize flow regime transitions via the slope of the pressure drop versus flow rate curve (from the ISS data) demonstrated the dominance of capillarity at low flow rates and the dependence of the gas holdup on pre-flow conditions 13.
@@
One-dimensional (1D) two-fluid models (TFMs) have been developed to understand these experimental observations from a fundamental (continuum) point of view 14. Assuming the flow to be 1D, uniform in the cross-section, and at steady-state, the full TFM equations can be simplified so that the unknowns are the axial pressure drop, the liquid holdup, and the interphase drag force 15, 16, 17. Following this approach, Zhang et al. 15 performed experiments and developed a 1D TFM at low Bond number $Bo$ (the ratio of gravitational to surface tension forces), within which the trickle flow regime is dominant. They incorporated Ergun-type correlations in the three interphase drag forces (i.e., the liquid-solid $f_{ls}$, gas-liquid $f_{gl}$, and gas-solid $f_{gs}$ ones) with appropriate relative velocities 6, 7, 8. At low $Bo$, capillarity implies a pressure jump (drop) across a gas-liquid interface: $p_{c}=p_{g}-p_{l}\neq 0$, where $p$ is pressure and the subscripts “$c$”, “$g$”, and “$l$” stand for “capillary,” “gas”, and “liquid,” respectively. For example, Zhang et al. 15 employed the correlation from Attou and Ferschneider 18 to model $p_{c}$ as a function of surface tension, particle diameter, gas holdup, and the ratio of gas and liquid densities. Next, estimating $dp_{c}/dx$ as $p_{c}/d_{p}$, where $p_{c}$ comes from the previous correlation, the respective pressures in the 1D gas and liquid phases’ momentum equations were eliminated to obtain a single momentum equation, with the liquid holdup being the only unknown. Finally, the pressure drop across the PBR was calculated by substituting the predicted liquid holdups, which were determined for different gas and liquid flow rates, obtaining a good agreement between the 1D TFM predictions and the experiments.
@@
Meanwhile, Salgi and Balakotaiah 16 used a modified Ergun correlation to calculate the pressure drop under microgravity conditions 10. They treated the liquid holdup and $f_{gl}$ as unknown in their model. They used a so-called “dynamic interaction term” from Motil et al. 10 to improve the fit to the experimental pressure drops. Further, a capillary pressure drop correlation, derived under terrestrial gravity conditions, was incorporated because there are no acceptable capillary pressure drop correlations for microgravity conditions. A linear stability analysis was performed to obtain $f_{gl}$ and liquid holdups at different gas and liquid flow rates. Similarly, Salgi and Balakotaiah 17 developed experimentally-based correlations for $f_{gl}$ from pulse velocity and frequency data to predict pressure drops through PBRs at both microgravity and terrestrial gravity conditions.
@@
In these prior 1D TFM studies, model calibration (of one or more fitting parameters) is key to accurately predicting pressure drops. Further, as Motil et al. 9 stated in their paper, “Ergun type expressions for $f_{gl}$ cannot be justified in the dispersed bubble flow regime.” However, it is the dispersed bubble and pulse flow regimes that are observed at microgravity conditions 9, 10, which we seek to model in this work. Therefore, the goal of the present work is to perform a data-driven calculation of the drag coefficient (fitting coefficient of our model) to obtain composite fits that yield a correlation for $f_{gl}$ at microgravity conditions.
@@
In addition to theoretical efforts, computational fluid dynamics (CFD) has been widely used to study resolved two-phase flows through PBRs. Specifically, an Euler-Euler TFM approach is used to predict pressure drop and liquid holdup within PBRs. Previous CFD studies19, 20, 21, 22, 23 considered an “empty” channel and incorporated the relevant closure relations for interphase drag terms (such as $f_{gl}$, $f_{ls}$, and $f_{gs}$) as momentum source terms, which implicitly accounts for the porous medium. In this modeling approach, one does not have to resolve the complex flow domain of interstitial void spaces to extract relevant flow parameters 24. In the present study, we follow this approach and incorporate the proposed $f_{gl}$ closure relations (developed from the 1D TFM model) as momentum source terms in CFD simulation. Then, we compare the pressure drop from these CFD simulations with the experimental pressure drop measured by NASA’s Packed Bed Reactor Experiment (PBRE), which accessed only the bubble and pulse flow regimes at microgravity, provided as data in NASA’s Physical Sciences Informatics (PSI) system 25. This comparison serves as an example demonstration of the utility of the proposed $f_{gl}$ correlations.
@@ 2 Mathematical model
Our specific aim is to obtain a closure relation for interphase gas-liquid drag force $f_{gl}$ in a TFM for flow through a PBR, based on the PBRE microgravity experiments 25. The idea of a TFM is to treat the carrier and dispersed phases of a two-phase flow as interpenetrating continua, each having its own thermophysical properties and obeying suitable balance laws 14. The two continua “interact” through the interphase drag force, denoted $f$. Therefore, the interphase drag force couples the momentum equations of the two phases. Throughout, it should be understood that $f$ usually represents a body force density (i.e., a force per unit volume), but we refer to it as “force” for clarity and simplicity.
In this section, we present our proposed procedure for obtaining $f_{gl}$. Specifically, in Sec. 2.1, we start from the basic mass and momentum balances under the TFM. In Sec. 2.2, we employ experimental data from NASA’s PBRE 25 to perform a data-driven inference and calibration of a model for $f_{gl}$. In Sec. 2.3, we demonstrate the self-consistency of the modeling approach by comparing the predicted pressure drops with the original experimental data.
@@ 2.1 Governing equations
The governing equations of gas-liquid flow through a PBR are those of conservation of mass and conservation of momentum. Following the predominant approach in the existing literature 16, 17, 15, and in the absence of detailed measurements of the interstitial flow and interface dynamics, the flow is considered predominantly 1D, with only the gradients along the axial (i.e., flow) direction of PBR being of interest in predicting the pressure drop. Under this assumption, the conservation of mass for liquid and gas phases, respectively, can then be simplified to:
@@
In a 1D TFM model of two-phase flow through a PBR, it is generally assumed (an approximation is made) that the mesoscale velocities are $v_{g}=v_{gs}/(\epsilon\phi_{g})$ and $v_{l}=v_{ls}/(\epsilon\phi_{l})$, where $v_{gs}$ and $v_{ls}$ are the corresponding gas or liquid superficial velocities, and $\epsilon$ is the PBR void fraction (porosity) 16, 17, 15. Next, the 1D conservation of momentum equations for the liquid and gas phases (based on the mesoscale velocities), respectively, are
@@
To close the TFM (Eqs. (1) and (2)), we need to specify the interphase momentum transfer terms, which are represented by the forces $f_{g}$ and $f_{l}$. Note that the expressions for $f_{g}$ and $f_{l}$ can each have a number of additive contributions, depending on the physics at hand. To arrive at expressions for $f_{g}$ and $f_{l}$, we assume that the interaction between the gas phase and the solid packing is negligible, i.e., $f_{gs}=0$, because the liquid wets the solid packing 16, 17. Then, we write the phase-specific forces in terms of the interphase ones as $f_{l}=f_{gl}-f_{ls}$ and $f_{g}=-f_{gl}$. We remind the reader that $f_{ls}$ is the interphase drag force between the liquid and solid phases, and $f_{gl}$ is the interphase drag force between the gas and liquid phases.
Next, it is standard to assume fully developed flow at steady state 16, 17, 15 and that the pressure gradients within the gas and liquid phases are equal 7, 8, 6. Incorporating these assumptions into Eqs. (2), we have
@@ 2.2.1 Two-phase Ergun correlation for the liquid-solid drag
Previous studies16, 17 used an Ergun-type 26, 27 equation to determine $f_{ls}$. An Ergun-type equation consists of additive viscous and inertial contributions to the pressure gradient across a PBR, but the relation can also be used to express the drag. Specifically, Salgi and Balakotaiah 16, Salgi and Balakotaiah 17 adapted the Ergun-type equation:
@@
where $d_{p}$ is the diameter of the spherical particles comprising the packing. Here, $f_{ls}=A_{ls}$ is a single-phase Ergun-type 26, 27 equation for the liquid phase, which has been adjusted by the denominator $\phi_{l}^{2}$ to account for the presence of a gas phase reducing the area available for the flow of the liquid phase. In the present study, we use Eq. (4) to model the drag between the liquid and solid phases where the coefficients $180$ and $1.8$ for vicious and inertia components respectively are from Motil et al. 9.
@@ 2.2.2 Data-driven calibration of the gas-liquid drag
Next, describe our approach to obtaining the closure relation (i.e., correlation) for $f_{gl}$ from the experimental pressure drop data obtained from NASA’s PBRE, which we accessed via NASA’s PSI 25. As indicated in the works of Salgi and Balakotaiah 16, bubble and pulse flow regimes are predominant at microgravity conditions, and the trickle flow regime does not occur. Hence, one cannot use an Ergun-type correlation for $f_{gl}$, with a relative velocity because that is valid only in the trickle flow regime 6, 7, 8. To circumvent this issue, we use a traditional drag force formulation: $f_{gl}=\beta(v_{g}-v_{l})$, where $\beta$ is the drag coefficient, and $v_{g}-v_{l}$ is the relative velocity. Using this formulation, the shared pressure gradient can be eliminated between Eqs. (3a) and (3b), while $f_{ls}$ is eliminated via Eq. (4), to obtain a single governing equation:
@@
Salgi and Balakotaiah 16 noted the absence of an acceptable correlation for the capillary pressure $p_{c}$ at microgravity conditions. While previous work 15 estimated $dp_{c}/dx$ as $p_{c}/d_{p}$ using a correlation based on terrestrial gravity conditions for $p_{c}$, we follow the reasoning of Attou and Ferschneider 18, namely “for a fully established gas-liquid flow, the gradient of the capillary pressure can be reasonably neglected compared with the gradient of pressure in each fluid,” justifying the shared pressure gradient assumption that led to Eq. (5).
@@
To quantify this dependence, we performed a set of linear curve fits on the $f_{gl}$ versus $Re^{*}_{gs}$ data at each $Re^{*}_{ls}$ to obtain correlations of the form: $f_{gl}=a+bRe^{*}_{gs}$, where $a$ and $b$ are, respectively, the intercept and slope of the curvefit at fixed value of $Re^{*}_{ls}$. To combine these multiple correlations into a single correlation, we perform a composite linear fit to the coefficients $a$ and $b$ (obtained across all $Re^{*}_{ls}$ values) with respect to $Re^{*2}_{ls}/Su_{l}$, where $Su_{l}=\rho_{l}d_{p}\sigma/\mu_{l}^{2}$ is the Suratman number, as shown in Fig. 1(b) and Fig. 1(c) of the forms: $a(\xi)=\alpha+\beta\xi$ and $b(\xi)=\gamma+\delta\xi$. We introduced $Su_{l}$ (a combined measure of the relative importance of surface tension to both inertial and viscous forces in the flow) here to implicitly account for surface tension (capillarity effects) between the gas and liquid phases, even if the capillary pressure gradient can be neglected in fully developed steady flow 18. Further, following the ideas leading Motil et al. 10 to their dynamic interaction term, we correlate $a$ and $b$ as functions of $Re^{*2}_{ls}/Su_{l}\sim We_{l}$ (the Weber number). Figure 1(a) shows that for largest values of $Re^{*}_{ls}$, the dependence of $f_{gl}$ on $Re^{*}_{gs}$ tends to be nonlinear for small values of $Re^{*}_{gs}$. However, since we perform a curve fit over the full range of data, these outliers do not significantly affect the trend and are not handled differently (see also the discussion of Fig. 3 below). In principle, fitting functions with more parameters could be devised to capture those points better, but given the uncertainties involved at low flow rates, such an approach would likely result in overfitting.
@@
$f_{gl}^{\text{Teflon}}$ $=291.6+4.22\times 10^{4}\frac{Re^{*2}_{ls}}{Su_{l}}+36.2Re^{*}_{gs}+4517\frac{Re^{*2}_{ls}Re^{*}_{gs}}{Su_{l}},$ (6a)
$f_{gl}^{\text{glass}}$ $=330.8+9.92\times 10^{4}\frac{Re^{*2}_{ls}}{Su_{l}}+27.2Re^{*}_{gs}+1.466\times 10^{4}\frac{Re^{*2}_{ls}Re^{*}_{gs}}{Su_{l}}.$ (6b)
Finally, if one were to justify that the capillary pressure gradient in a two-phase flow in microgravity conditions is not negligible, it can be added back into Eq. (5), potentially via a correlation of the form $p_{c}/d_{p}$ 15. In such a case, the fitting process can be repeated, and the curve-fit coefficients would change but the methodology remains the same. It must also be noted that the proposed correlation for $f_{gl}$ is applicable for monodisperse Teflon and glass packing materials and cannot be generalized for other packing materials with different sizes. This drawback stems from the limited nature of the experimental data from NASA’s PBRE 25. Furthermore, the proposed correlations can be used to predict bulk properties, such as the gas-liquid pressure drop across a PBR at microgravity conditions, but not interstitial properties, such as the flow regime, which are limited to bubble and pulse flow in NASA’s PBRE at microgravity conditions.
Figure 3: Verification of the proposed $f_{gl}$ correlations in Eq. (6) (and hence the approach) by comparing the model and experiment’s $-dp/dz$ variation with $Re^{*}_{gs}$ at different $Re^{*}_{ls}$ for (a) Teflon packing and (b) glass packing PBRE data. As expected, the model predictions ($\bigtriangleup$) at discrete data points coincide exactly with the values from experiments ($\times$). The arrow indicates the direction of increasing $Re^{*}_{ls}$. The dashed lines are a guide to the eye and do not represent any trends. Note the logarithmic abscissa.